OpenStudy (anonymous):

Determine the zeros of f(x) = x^3 – 3x^2 – 16x + 48

OpenStudy (anonymous):

factor by grouping and set to zero

OpenStudy (agreene):

x= -4,3,4

OpenStudy (anonymous):

x^3 – 3x^2 – 16x + 48 = 0 x^2(x-3) -16(x-3) = 0 (x^2-16)(x-3) = 0 x^2 = 16; or x = 3 x = +4, -4 or 3